# Incremental Generation is Necessity and Sufficient for Universality in Flow-Based Modelling

### v1

@@ Abstract
Using a new topological-dynamical argument, we first prove an impossibility theorem: the class of all single-step autonomous flows, independently of the architecture, width, depth, or Lipschitz activation of the underlying neural network, is meagre and therefore not universal in the space of orientation-preserving homeomorphisms of $0,1]^{d}$. By exploiting algebraic properties of autonomous flows, we conversely show that every orientation-preserving Lipschitz homeomorphism on $0,1]^{d}$ can be approximated at rate $\mathcal{O}(n^{-1/d})$ by a composition of at most $K_{d}$ such flows, where $K_{d}$ depends only on the dimension. Under additional smoothness assumptions, the approximation rate can be made dimension-free, and $K_{d}$ can be chosen uniformly over the class being approximated. Finally, by linearly lifting the domain into one higher dimension, we obtain structured universal approximation results for continuous functions and for probability measures on $0,1]^{d}$, the latter realized as pushforwards of empirical measures with vanishing $1$-Wasserstein error.
@@ 1 Introduction
$\sup_{\omega}\,\|\varphi(Z(\omega))-\psi(Z(\omega))\|\leq d_{\infty}(\phi,\psi)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sup_{x\in\mathbb{R}^{d}}\,\|\varphi(x)-\psi(x)\|$ (2)
@@
Unfortunately, $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is not a vector space which places it outside the scope of the available constructive approximation toolbox pinkus2012n; CABook2; cohen2022optimal; petrova2023Lipschitz, which assumes a Banach space structure—our concept class $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ possesses a rich and well-studied infinite-dimensional topological group structure mather1974commutators; thurston1974foliations; fathi1980structure; mather1984curious; banyaga1997structure; haller2013smooth; fukui2019uniform that we will exploit extensively. Moreover, unlike signature-based machine learning methods, cf. gyurko2013extractingsignature; chevyrev2016primer; cuchiero2023signaturecalibration; hamida2005signaturetest; cass2024lecturenotes; bayer2025optimalstopping, rooted in rough path theory lyons1998differential; gubinelli2004controlling; HamblyLyons2010, we do not have access to a global exponential map (in the sense of infinite-dimensional Lie groups; cf. Kriegl1997; michor2023regularity) to transfer results from linear spaces onto $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$, as in kratsios2020non; kratsios2022universalpapon. In short, there are currently no available tool to apply classical approximation results from linear spaces onto $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ using global exponential map, we are required to develop new approximation-theoretic techniques suited to its non-vectorial structure, where our approximators/hypothesis class consists only of (neural network-based) homeomorphisms in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$. Our approximators preserve the structure of the maps in (1), whereas replacing $\varphi$ with an arbitrary, e.g. non-invertible neural network as in the standard GAN framework radford2016unsupervised; arjovsky2017wasserstein; peyre2019computational; rout2022generative; korotin2021wasserstein; korotin2023neural; kolesov2024energy would prevent the use of denoising-based training.
@@
$\varphi$ $=\varphi_{G}\circ\dots\circ\varphi_{1},$ (3)
for some $G\in\mathbb{N}_{+}$, where for each $l\in\{1,\dots,G\}$, the map $\varphi_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a time-$1$ flow of an autonomous ODE; that is, $\varphi_{l}$ sends any $x\in\mathbb{R}^{d}$ to the time-$1$ solution $x_{1}^{x}$ of the autonomous ODE with initial condition $x$
@@
where $V_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a Lipschitz vector field, parameterized by a ReLU multilayer perceptron (MLP)whose weights matrices has finite operator norm. Under this formalism, a flow-based generative model $\varphi$, as in (3), is called incremental if $G>1$ and it is called non-incremental if $G=1$. We note some $L^{p}$-version, for finite $p$, of these results in this direction are known via controllability arguments ruiz2024control, however these results come at a cost as they require the user to have active control of the vector field of the Neural ODE (i.e. non-autonomous or controlled NeuralODEs) which is effectively the case where $G=\infty$.
Figure 1: Visualizing Incremental Flow-Based Generation (3): In a denoising-type, flow-based incremental generator, an image $x$ (left) is mapped to noise by the invertible sequence $\varphi_{G}^{-1}\circ\cdots\circ\varphi_{1}^{-1}$. The model is trained to undo this via $\varphi_{1}\circ\cdots\circ\varphi_{G}$. At inference, a noise sample $Z$ is injected and propagated through $\varphi_{1},\dots,\varphi_{G}$ to synthesize an image  [^3] Image credit: Sans Soleil (1983), written and directed by Chris Marker..
Here, $G=2$. Each arrow depicts the vector field—visualized in the panel directly below—that induces one step of the incremental flow. By contrast, a non-incremental generator attempts a single-shot mapping from left to right, while non-denoising pipelines (e.g., GANs goodfellow2014generative) do not enforce invertibility.
In contrasts, incremental flow-based generative models can be approximately represented as a single non-autonomous flow, i.e., where the vector field in (4) is allowed to be time-dependent but undergoes only finitely many changes in direction; which directly allows the conversion of our theory of incremental flow-based generation to time-dependent generation. Such time-dependent models differ from controlled neural ODEs kidger2020neural; cuchiero2020deep; cirone2023neural, which mimic controlled rough paths morrill2021neural; pmlr-v235-walker24a and can potentially adapt their vector fields arbitrarily often over time resulting in a higher parametric complexity. The following question drives our manuscript and an affirmative answer to it would imply a concrete mathematical support for the advantage of incrimental generation over classical non-incremental generation to support the undeniable success of incremental generation
@@ 1.1 Main Construction
Our first result (Theorem 1) shows that incremental generation, i.e. when $G>1$, is necessary for universality of flow-based generative models in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ for $d\in\mathbb{N}_{+}$ and $d>1$. Specifically, there exists a $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ that cannot be approximated by the flow of any autonomous neural ODE, regardless of network depth, width, or choice of activation function. We deduce this from a more general result of ours showing that: the class of all flows of autonomous ODEs is meagre (in particular, not dense) in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$. This result extends the differential topology results of Varandas_2020__Meag to the non-compact case and the classical embedding results of fort1955embedding; jones1972embedding; palis1974vector to the non-smooth, non-compact setting.
@@
Our final main result (Theorem 9) shows, somewhat surprisingly, that $G$ never needs to become arbitrarily large when approximating diffeomorphisms. Even more strikingly, we prove that there exists a dimensional constant $K_{d}>0$ such that one can constrain $G\leq K_{d}$, independantly of the target diffeomorphism being approximated and maintain universality in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
@@ 2.1.1 Homeomorphisms, Flows and Flowability
This section contains all necessary terminology and background needed to formulate our main results. Additional background required only for proofs is contained in Appendix C. We say a function is supported on a set $S$ if it is zero out of that set and a function is supported if it is the identity out of that set. For practical reasons said in the introduction section, we need to work with compactly supported/supported functions and vector fields; meaning that $S$ is contained in a compact set.
@@ 2.1.2 From Neural ODEs to Incremental Flow-Based Generators
Let $\Delta\in\mathbb{N}_{+}$ and consider a multi-index $\mathbf{d}\ \stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\ [d_{1},\ldots,d_{\Delta+1}]\in\mathbb{N}_{+}^{\Delta+1}$. The class $\mathcal{NN}(\mathbf{d})$ consists of all multilayer perceptrons with $\sigma\in C(\mathbb{R})$ Lipschitz-activation function ($\sigma$-MLPs)
@@
Here, for $l=1,\ldots,\Delta$, $\mathbf{W}^{(l)}$ is a $d_{l+1}\times d_{l}$ matrix and $\mathbf{b}^{(l)}\in\mathbb{R}^{d_{l+1}}$, and $\sigma\ \bullet$ denotes componentwise application of the $\sigma$ function.
Fix a Lipschitz activation function $\sigma\in C(\mathbb{R})$ and real numbers $a\leq b$. A $\sigma$-neural ODE of respective depth and width $\Delta,W\in\mathbb{N}_{+}$ is a flowable homeomorphism $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ for which there exists a $\sigma$-MLP $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ of depth $\Delta$ and width $W$ supported in $a,b]^{d}$ such that $\varphi$ is time one solution of (4) with vector field $\Phi$. The class of all $\sigma$-neural ODEs supported on $a,b]^{d}$ is denoted as $\mathrm{NODE}_{\sigma}([a,b]^{d})$. The class of incrimental flow-based generators with activation function $\sigma$ denoted as $\mathrm{IFG}_{\sigma}([a,b]^{d})$ is defined as all functions of $\operatorname{\mathcal{H}}_{d}([a,b]^{d})$ that can be written as composition of at least two or more but finitely many $\sigma$-neural ODEs
@@ 2.2 Notation
  • Denote (open Euclidean) ball with radious r around point p as $B_{r}(p)=\{x\in\mathbb{R}^{d}|\ \|x-p\|<r\}$ where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
@@
  • For $s\in\mathbb{N}\cup\{\infty\}$ let $\mathcal{X}^{s}$ be the set of $C^{s}$ vector fields on $\mathbb{R}^{d}$ supported on $0,1]^{d}$. when $s=0$, we only consider Lipschitz continuous vector fields for $\mathcal{X}^{0}$.
@@
  • Fix a granularity $G\in\mathbb{N}_{+}$ and a smoothness $s\in\mathbb{N}$. A homeomorphism $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ belongs to $\operatorname{\mathcal{H}}_{d}^{G,s}([0,1]^{d})$ if and only if:
@@
      Representation: There exist vector fields $V_{1},\dots,V_{G}\in\mathcal{X}^{s}$ such that
      $\varphi=\bigcirc_{g=1}^{G}\,\operatorname{Flow}{}(V_{g}).$ (7)
@@
      Minimality: There is no integer $1\leq\tilde{G}<G$ and vector fields $\tilde{V}_{1},\dots,\tilde{V}_{G}\in\mathcal{X}^{s}$ such that
      $\varphi=\bigcirc_{g=1}^{\tilde{G}}\,\operatorname{Flow}{}(\tilde{V}_{g}).$
@@
  • Given $E\subseteq\mathbb{R}^{d}$, $d$-tuple $\alpha=[\alpha_{1},\alpha_{2},\cdots,\alpha_{d}]^{T}\in\mathbb{N}^{d}$ and functions $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ let :

### v2

# Incremental Generation is Necessary and Sufficient for Universality in Flow-Based Modelling

@@ Abstract
Using a new topological-dynamical argument, we first prove an impossibility theorem: the class of all single-step autonomous flows, independently of the architecture, width, depth, or Lipschitz activation of the underlying neural network, is not universal in the space of orientation-preserving homeomorphisms of $0,1]^{d}$. By exploiting algebraic properties of autonomous flows, we conversely show that every orientation-preserving Lipschitz homeomorphism on $0,1]^{d}$ can be approximated at rate $\mathcal{O}(n^{-1/d})$ by a composition of at most $T_{d}$ such flows, where $T_{d}$ depends only on the dimension. Under additional smoothness assumptions, the approximation rate can be made dimension-free, and $T_{d}$ can be chosen uniformly over the class being approximated. Finally, by linearly lifting the domain into one higher dimension, we obtain structured universal approximation results for continuous functions and for probability measures on $0,1]^{d}$, the latter realized as pushforwards of empirical measures with vanishing $1$-Wasserstein error.
@@ 1 Introduction
where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{D}$. Generation is performed by sampling $Z$, which induces a sample of $X$. The transformation $\varphi$ is required to be continuous, often Lipschitz, to ensure numerical stability during generation gulrajani2017improved; cisse2017parseval; miyato2018spectral. We consider denoising-based generative models, such as normalizing flows rezende2015variational; kingma2018glow which conceptually build on their variational autoencoder (VAE) predecessors kingma2013auto. These models require the transformation $\varphi$ to be continuously invertible and leverage the inverse map $\varphi^{-1}:\varphi(\mathbb{R}^{d})\to\mathbb{R}^{d}$ to train the model: given i.i.d. samples from $X$, the model learns to map them back to the noise source $Z$. The invertibility, thus injectivity of $\varphi$ necessitates that $D\geq d$; we focus on the difficult critical case where $D=d$ and $\varphi$ is surjective (other cases are a simple consequence which we discussed below). Lastly, keeping with the approximation-theoretic tradition in deep learning mhaskar2016deep; yarotsky2017error; petersen2018optimal; elbrachter2021deep; kratsios2022do; zhang2022deep; siegel2023optimal, we will focus only on maps $\varphi$ that are supported (in the sense of differential topology, cf. (Banyaga_1997__DiffeoBook, Chapter 2)) on the unit hypercube $0,1]^{d}$; that is, $\varphi(x)=x$ for every $x\notin[0,1]^{d}$. Consequently, Alexander’s trick implies that every such homeomorphism must be orientation preserving. This means that, the largest concept class of denoising-type generative models consists of all random variables $X$, as in (1), for which the transformation $\varphi$ belongs to the class $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ of orientation-preserving homeomorphism supported on $0,1]^{d}$. Whenever both $X$ and $Z$ have a Lebesgue density on $0,1]^{d}$ then the existence of such a $\varphi$ is guaranteed by the main result of dacorogna1990partial; regardless of $Z$ and $X$.
@@
$\sup_{\omega}\,\|\varphi(Z(\omega))-\psi(Z(\omega))\|\leq d_{\infty}(\varphi,\psi)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sup_{x\in\mathbb{R}^{d}}\,\|\varphi(x)-\psi(x)\|$ (2)
@@
Unfortunately, $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is not a vector space which places it outside the scope of the available constructive approximation toolbox pinkus2012n; CABook2; cohen2022optimal; petrova2023Lipschitz, which assumes a Banach space structure—our concept class $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ possesses a rich and well-studied infinite-dimensional topological group structure mather1974commutators; thurston1974foliations; fathi1980structure; mather1984curious; banyaga1997structure; haller2013smooth; fukui2019uniform that we will exploit extensively. Moreover, unlike signature-based machine learning methods, cf. gyurko2013extractingsignature; chevyrev2016primer; cuchiero2023signaturecalibration; hamida2005signaturetest; cass2024lecturenotes; bayer2025optimalstopping, rooted in rough path theory lyons1998differential; gubinelli2004controlling; HamblyLyons2010, we do not have access to a global exponential map (in the sense of infinite-dimensional Lie groups; cf. Kriegl1997; michor2023regularity) to transfer results from linear spaces onto $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$, as in kratsios2020non; kratsios2022universalpapon. In short, there are currently no available tools to apply classical approximation results from linear spaces onto $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ using global exponential map. we are required to develop new approximation-theoretic techniques suited to its non-vectorial structure, where our approximators/hypothesis class consists only of (neural network-based) homeomorphisms in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$. Our approximators preserve the structure of the maps in (1), whereas replacing $\varphi$ with an arbitrary, e.g. non-invertible neural network as in the standard GAN framework radford2016unsupervised; arjovsky2017wasserstein; peyre2019computational; rout2022generative; korotin2021wasserstein; korotin2023neural; kolesov2024energy prevents the use of denoising-based training.
@@
$\varphi$ $=\varphi_{T}\circ\dots\circ\varphi_{1},$ (3)
for some $T\in\mathbb{N}_{+}$, where for each $l\in\{1,\dots,T\}$, the map $\varphi_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a time-$1$ flow of an autonomous ODE; that is, $\varphi_{l}$ sends any $x\in\mathbb{R}^{d}$ to the time-$1$ solution $x_{1}^{x}$ of the autonomous ODE with initial condition $x$
@@
where $V_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a Lipschitz vector field, parameterized by a ReLU multilayer perceptron (MLP)whose weights matrices has finite operator norm. Under this formalism, a flow-based generative model $\varphi$, as in (3), is called incremental if $T>1$ and it is called non-incremental if $T=1$. We note some $L^{p}$-version, for finite $p$, of these results in this direction are known via controllability arguments ruiz2024control; scagliotti2023deep; cuchiero2020deep; cheng2025interpolation, however these results come at a cost as they require the user to have active control of the vector field of the Neural ODE (i.e. non-autonomous or controlled Neural ODEs) which is effectively the case where $T=\infty$.
Figure 1: Visualizing Incremental Flow-Based Generation (3): In a denoising-type, flow-based incremental generator, an image $x$ (left) is mapped to noise by the invertible sequence $\varphi_{T}^{-1}\circ\cdots\circ\varphi_{1}^{-1}$. The model is trained to undo this via $\varphi_{1}\circ\cdots\circ\varphi_{T}$. At inference, a noise sample $Z$ is injected and propagated through $\varphi_{1},\dots,\varphi_{T}$ to synthesize an image  [^2] Image credit: Sans Soleil (1983), written and directed by Chris Marker..
Here, $T=2$. Each arrow depicts the vector field—visualized in the panel directly below—that induces one step of the incremental flow. By contrast, a non-incremental generator attempts a single-shot mapping from left to right, while non-denoising pipelines (e.g., GANs goodfellow2014generative) do not enforce invertibility.
In contrasts, incremental flow-based generative models can be approximately represented as a single non-autonomous flow, i.e., where the vector field in (4) is allowed to be time-dependent but undergoes only finitely many changes in direction; which directly allows the conversion of our theory of incremental flow-based generation to time-dependent generation. Such time-dependent models differ from controlled neural ODEs kidger2020neural; cuchiero2020deep; cirone2023neural, which mimic controlled rough paths morrill2021neural; pmlr-v235-walker24a and can potentially adapt their vector fields arbitrarily often over time resulting in a higher parametric complexity. The following question drives our manuscript and an affirmative answer to it would imply a concrete mathematical support for the advantage of incremental generation over classical non-incremental generation to support the undeniable success of incremental generation
@@ 1.1 Main Construction
Our first result (Theorem 2) shows that incremental generation, i.e. when $T>1$, is necessary for universality of flow-based generative models in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ for $d\in\mathbb{N}_{+}$ and $d>1$. Specifically, there exists a $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ that cannot be approximated by the flow of any autonomous neural ODE, regardless of network depth, width, or choice of activation function. We deduce this from a more general result of ours showing that: the class of all flows of autonomous ODEs is non-universal (in particular, not dense) in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$. This result extends the differential topology results of Varandas_2020__Meag to the non-compact case and the classical embedding results of fort1955embedding; jones1972embedding; palis1974vector to the non-smooth, non-compact setting.
@@
Our final main result (Theorem 10) shows, somewhat surprisingly, that $T$ never needs to become arbitrarily large when approximating diffeomorphisms. Even more strikingly, we prove that there exists a dimensional constant $T_{d}>0$ such that one can constrain $T\leq T_{d}$, independently of the target diffeomorphism being approximated and maintain universality in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
@@ 2.1.1 Homeomorphisms, Flows and Flowability
This section contains all necessary terminology and background needed to formulate our main results. Additional background required only for proofs is contained in Appendix D. We say a function is supported on a set $S$ if it is zero out of that set and a function is supported if it is the identity out of that set. For practical reasons said in the introduction section, we need to work with compactly supported/supported functions and vector fields; meaning that $S$ is contained in a compact set.
@@ 2.1.2 From Neural ODEs to Incremental Flow-Based Generators
Let $\Delta\in\mathbb{N}_{+}$ and consider a multi-index $\mathbf{d}\ \stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\ [d_{1},\ldots,d_{\Delta+1}]^{\top}\in\mathbb{N}_{+}^{\Delta+1}$. The class $\mathcal{NN}(\mathbf{d})$ consists of all multilayer perceptrons with $\sigma\in C(\mathbb{R})$ Lipschitz-activation function ($\sigma$-MLPs)
@@
Here, for $l=1,\ldots,\Delta$, $\mathbf{W}^{(l)}$ is a $d_{l+1}\times d_{l}$ matrix and $\mathbf{b}^{(l)}\in\mathbb{R}^{d_{l+1}}$, and $\sigma\ \bullet$ denotes componentwise application of the $\sigma$ function. Fix a Lipschitz activation function $\sigma\in C(\mathbb{R})$ and real numbers $a\leq b$. A $\sigma$-neural ODE of respective depth and width $\Delta,W\in\mathbb{N}_{+}$ is a flowable homeomorphism $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ for which there exists a $\sigma$-MLP $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ of depth $\Delta$ and width $W$ supported in $a,b]^{d}$ such that $\varphi$ is time one solution of (4) with vector field $\Phi$; also, the class of all $\sigma$-neural ODEs supported on a set $P\subseteq\mathbb{R}^{d}$ is denoted as $\mathrm{NODE}_{\sigma}(P)$. The class of incremental flow-based generators with activation function $\sigma$ denoted as $\mathrm{IFG}_{\sigma}(P)$ is defined as all functions of $\operatorname{\mathcal{H}}_{d}(P)$ that can be written as composition of at least two or more but finitely many $\sigma$-neural ODEs.
@@ 2.2 Notation
  • Denote (open Euclidean) ball with radius r around point p as $B_{r}(p)=\{x\in\mathbb{R}^{d}|\ \|x-p\|<r\}$ where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
@@
  • For $s\in\mathbb{N}\cup\{\infty\}$ let $\mathcal{X}^{s}$ be the set of $C^{s}$ vector fields on $\mathbb{R}^{d}$ supported on $0,1]^{d}$. Moreover, we consider $\mathcal{X}^{0}$ to be the set of Lipschitz continuous vector fields.
@@
  • Fix an Incrementality $T\in\mathbb{N}_{+}$ and a smoothness $s\in\mathbb{N}$. A homeomorphism $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ belongs to $\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ if and only if:
@@
      Representation: There exist vector fields $V_{1},\dots,V_{T}\in\mathcal{X}^{s}$ such that
      $\varphi=\bigcirc_{t=1}^{T}\,\operatorname{Flow}{}(V_{t}).$ (7)
@@
      Minimality: There is no integer $1\leq\tilde{T}<T$ and vector fields $\tilde{V}_{1},\dots,\tilde{V}_{T}\in\mathcal{X}^{s}$ such that
      $\varphi=\bigcirc_{t=1}^{\tilde{T}}\,\operatorname{Flow}{}(\tilde{V}_{t}).$
@@
  • Given $E\subseteq\mathbb{R}^{d}$, $d$-tuple $\alpha=[\alpha_{1},\alpha_{2},\cdots,\alpha_{d}]^{\top}\in\mathbb{N}^{d}$ and functions $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ let :